Calculate the pH after 0.15 mole of NaOH is added to 1.07 L of a solution that is 0.59 M HNO2 and 1.05 M KNO2, and calculate the pH after 0.30 mole of HCl is added to 1.07 L of the same solution of HNO2 and KNO2.

0.15 mole of NaOH __________

0.30 mole of HCl __________
Answer #1

First calculate no of moles of HNO2 and KNO2 in buffer solution

We have , Molarity = No of moles of solute / Volume of solution in L

Therefore, No of moles of Solute = Molarity x Volume of solution in L

No of moles of HNO2= 0.59 mol / L x 1.07 L

= 0.63 mol

No of moles of KNO2 = 1.05 mol / L x 1..07 L

= 1.12 mol

pH of buffer after 0.15 mol NaOH

When NaOH is added to a buffer it will react with acidic component of buffer namely HNO2 and produces NO2-base  . Due to addition of NaOH , [ HNO2 ] decreases while that of [NO2-] increases. Hence new concentration will be

Concentration HNO2 NaOH NO2-
Initial 0.63 0.15 1.12
Change - 0.15 -0.15 +0.15
Equilibrium 0.48 0.00 1.27

We know that, pH of buffer is calculated by using Henderson's equation

pH = pKa + log [ salt] / [Acid]

Here pKa of HNO2 = 3.15 , [ salt] = [ NO2-] = 1.27 and [Acid] = [ HNO2 ] = 0.48

Therefore, pH = 3.15 + log 1.27 / 0.48 ( Here volume is not considered as it is same for both acid and salt)

= 3.15 + 0.42

= 3.57

ANSWER: pH of buffer after addition of 0.15 mol NaOH = 3.57

When HCl is added to a buffer it will react with basic component of buffer namely NO2- and produces HNO2  acid . Due to addition of HCl , [ HNO2 ] increases while that of [NO2-] decreases. Hence new concentration will be

Concentration HNO2 HCl NO2-
Initial 0.63 0.30 1.12
Change + 0.30 -0.30 -0.30
Equilibrium 0.93 0.00 0.82

We know that, pH of buffer is calculated by using Henderson's equation

pH = pKa + log [ salt] / [Acid]

Here pKa of HNO2 = 3.15 , [ salt] = [ NO2-] = 0.82 and [Acid] = [ HNO2 ] = 0.93

Therefore, pH = 3.15 + log 0.82 / 0.93 ( Here volume is not considered as it is same for both acid and salt)

= 3.15 - 0.055

= 3.10

ANSWER: pH of buffer after addition of 0.30 mol HCl = 3.10